# Debiased machine learning for combining probability and non-probability survey data

### v1

@@ 6 Ratio estimator
This estimator $\hat{\theta}_{{\rm CLW}2}$ may be more efficient than $\hat{\theta}_{1}$, for the same reason that the Hajek estimator may be more efficient than the Horwitz-Thompson estimator. In CLW’s simulation study with parametric nuisance models, $\hat{\theta}_{{\rm CLW}2}$ did indeed have smaller variance than $\hat{\theta}_{1}$.
@@
The mean-zero term $\sqrt{M}n^{-1}\sum_{i=1}^{n}\{m_{0}(X_{i})-Y_{i}\}$ in expression (7) is $O_{p}(1)$ and does not vanish asymptotically. However, when the population is large compared to Samples A and B, the variance of this term is small compared to the variance of the other terms in expression (7). We see that if this term is ignored, equation (7) is the same as equation (4) but with $m_{0}(X_{i})$ in the first term replaced by $m_{0}(X_{i})-n^{-1}\sum_{i^{\prime}=1}^{n}m_{0}(X_{i^{\prime}})$.
Ignoring the term $\sqrt{M}n^{-1}\sum_{i=1}^{n}\{m_{0}(X_{i})-Y_{i}\}$ in equation (7), the asymptotic variance of $\hat{\theta}_{{\rm CLW}2}-\bar{Y}$ (and hence of $\hat{\theta}_{2}-\bar{Y}$) is given by CLW, along with an estimator of this variance. This is described in more detail by Seaman et al. (2025)[25] (specifically, the formulae for $\mu_{\rm DR2}$ when both nuisance models are correctly specified).
@@ 7 Targeted maximum likelihood
A popular alternative to estimator $\hat{\theta}_{\rm iid}$ for iid data is the targeted maximum likelihood estimator (TMLE). The TMLE method involves modifying initial estimate $\hat{m}_{\rm iid}$ of $E(Y\mid X)$ to $\hat{m}^{*}_{\rm iid}$ in such a way that the simple regression imputation estimator $n^{-1}\sum_{i=1}^{n}\hat{m}_{\rm iid}^{*}(X_{i})$ of $E(Y)$ can be used. We now describe an analogous TMLE method for our survey sample situation.
@@
for some $\epsilon_{k}$. Alternatively, if $Y$ is bounded by zero and one, i.e. $P(0<Y<1)=1$, then $\hat{m}_{k}(x;\epsilon_{k})$ can be defined by equation (8) or equation (9):
@@ 8 Calculating $\hat{\pi}^{B}_{k}$ and $\hat{m}_{k}$
So far, we have been agnostic about how to calculate $\hat{\pi}^{B}_{k}$ and $\hat{m}_{k}$, except to specify that they use, respectively, $(X,R^{A},R^{B})$ values and $(X,Y)$ values of individuals not in fold $k$. Estimating $m_{0}(X)=E(Y\mid X)$ is relatively straightforward, because $Y_{1},\ldots,Y_{n}$ are assumed to be independent given $X_{1},\ldots,X_{n}$. This is a standard problem for which many data-adaptive methods could be used.
@@
This is the log likelihood for a binomial regression model of $R^{B}$ on $X$ in the population of $n$ individuals, using the individuals in Sample A weighted by their inverse sampling probabilities to represent the population, so that the ‘failures’ or ‘non-events’ are represented by the weighted Sample A minus the individuals in Sample B. Ferri-Garcia et al. (2022)[15] (see also [10]) use the pseudo-likelihood together with a classification and regression tree to estimate $\pi^{B}_{0}(X)$.
The approximate log pseudo-likelihood is the same as expression (11) but omitting the $R^{B}_{i}\{1-\pi^{B}(X_{i})\}$ term. When Sample B is small compared to the population, this term will be negligible compared to the other terms in expression (11). The approximate pseudo-likelihood has the advantage that it can be used quite generally with data-adaptive methods that predict a binary outcome and that allow for weights. We acknowledge CLW’s criticism of the approximate pseudo-likelihood in situations where Sample B is not a small fraction of the population, but we focus in Section 9 on scenarios where Sample B is small compared to the population, a situation which would be quite common in practice.
@@ 9 Simulation study
We simulated a finite population with $J=1000$ clusters. Each cluster $j$ contains $H_{jq}$ households of $q$ individuals ($q=1,2,3$), where $H_{jq}$ is negatively binomially distributed with mean 100 and variance 400. Hence, the expected number of individuals in a cluster is 600, and the expected population size is 600,000. For each individual $i$ in cluster $j$, continuous variables $X_{1i}$ and $X_{2i}$ and binary variables $X_{3i}$ and $X_{4i}$ were generated independently, with expectations that depend on the total number of households in the cluster $j$, the size $q$ of the household to which individual $i$ belongs, and a cluster-level random effect. This fixed population was used for all the simulations.
For each simulated dataset, $Y_{1},\ldots,Y_{n}$ were generated independently from a normal distribution with mean that is a function of $X_{i}=(X_{1i},X_{2i},X_{3i},X_{4i})$ and with variance 1. Sample A was drawn by sampling $M$ clusters with replacement from the $J$ clusters using Sampford sampling, with the probability of sampling cluster $j$ being proportional to the total number of households in cluster $j$. From each of these $M$ sampled clusters, $n_{\rm house}$ households were sampled using simple random sampling without replacement. Then one individual was sampled at random from each of these $n_{\rm house}$ sampled households. Thus, Sample A contained $M\times n_{\rm house}$ individuals. The probability of sampling individual $i$ for Sample B was a function of $X_{i}$.
We considered six scenarios. In Scenarios 1 and 2, $m_{0}(X)$ is linear in $X$ and $\pi^{B}_{0}(X)$ follows a logistic regression model with only main effects for $X_{1}$, $X_{2}$, $X_{3}$ and $X_{4}$. Specifically, $m_{0}(X)=X_{1}+X_{2}+2X_{3}+X_{4}$ and $\mbox{logit }\pi^{B}_{0}(X)=\alpha_{\rm int}+0.5X_{1}+X_{2}+0.5X_{3}+X_{4}$. Both these scenarios use $M=150$ and $n_{\rm house}=20$. They differ in the value of $\alpha_{\rm int}$, and hence the expected size of Sample B. In Scenario 1, $\alpha_{\rm int}=-6.2$ and the expected size of Sample B is 7000; in Scenario 2, $\alpha_{\rm int}=-7.5$ and the expected size of Sample B is 2000. In Scenarios 3–6, $m_{0}(X)$ is a non-linear function of $X$ and the logistic regression model for $\pi^{B}_{0}(X)$ includes an interaction and a quadratic term. Specifically, $m_{0}(X)=0.5X_{1}+0.5X_{2}+2X_{3}+X_{4}+2X_{1}X_{3}+X_{2}^{2}$ and $\mbox{logit }\pi^{B}_{0}(X)=-6.4+0.25X_{1}+0.5X_{2}+0.5X_{3}+X_{4}+X_{1}X_{3}+0.5X_{2}^{2}$. The expected size of Sample B is 7500. These four scenarios differ in the values of $M$ and $n_{\rm house}$. Specifically, Scenarios 3–6 use, respectively, $(M,n_{\rm house})=(150,20)$, $(50,20)$, $(150,5)$ and $(50,5)$.
The following estimators were applied to each of 100 simulated datasets for each of the six scenarios.
@@
: $\hat{\theta}_{1}$ using Highly Adaptive LASSO (HAL) to estimate $\pi^{B}_{0}$ and $m_{0}$ with cross-fitting and 5 folds
@@
: $\hat{\theta}_{2}$ using HAL to estimate $\pi^{B}_{0}$ and $m_{0}$ with cross-fitting and 5 folds
@@
: $\hat{\theta}_{\rm TMLE1}$ using HAL to estimate $\pi^{B}_{0}$ and $m_{0}$ with cross-fitting and 5 folds
@@
: $\hat{\theta}_{\rm TMLE2}$ using HAL to estimate $\pi^{B}_{0}$ and $m_{0}$ with cross-fitting and 5 folds
Of these, DR1, DR2clw and KH were described by CLW. For the methods using HAL, we set $\delta=0.01$ and used $L=4$ sets of clusters based on quartiles of the distribution of $\pi^{C}_{j}$ (as described at the end of Section 5).
In addition, we calculated the estimators using generalised boosting models (GBM), in place of HAL, to estimate the nuisance models. These will be denoted ‘DR1.gbm5’ and so on. We also applied the estimators that used HAL or GBM without cross-fitting. These will be denoted ‘DR1.hal1’, ‘DR1.gbm1’, and so on.
Results are shown in Tables 1, 2, 3, 4, 5 and 6 for Scenarios 1–6, respectively.
In Scenarios 1 and 2, the methods using parametric nuisance models, viz. DR1, DR2clw, DR2, TMLE1, TMLE2 and KH, yield approximately unbiased point estimates and standard error estimates, and the coverage of 95% confidence intervals is correct. The estimators DR2clw and TMLE2 are more efficient than the DR1 and TMLE1 estimators, and DR2 is, as expected, as efficient as DR2clw. Using HAL to estimate the nuisance functions provides no benefit in these scenarios, where the parametric nuisance models are correctly specified, but also no loss of efficiency or coverage. Using GBM leads to some bias and a small deterioriation in coverage, but bias is still small and coverage not far from its nominal level.
In Scenarios 3–6, the methods using parametric nuisance models yield biased estimates and confidence intervals with poor coverage. When HAL is used, the bias is small and the coverage close to its nominal level. Results from GBM are similar, although the bias can be slightly greater than for HAL. Using cross-fitting with five folds appears to provide no performance improvement over not using cross-fitting in these scenarios but also no loss of efficiency.
bias empSE SEhat cover
HT 0.002 0.046 0.046 94
Haj 0.000 0.037 0.037 95
naive -0.288 0.035
DR1 0.001 0.044 0.044 94
DR2clw -0.000 0.035 0.035 95
DR2 -0.000 0.035 0.035 95
TMLE1 0.001 0.044 0.044 94
TMLE2 -0.000 0.035 0.035 95
KH 0.001 0.044 0.044 94
DR1.hal5 0.002 0.044 0.044 93
DR2.hal5 0.001 0.035 0.034 94
TMLE1.hal5 0.002 0.044 0.044 93
TMLE2.hal5 0.000 0.035 0.034 94
DR1.hal1 0.002 0.044 0.044 93
DR2.hal1 0.000 0.034 0.034 94
TMLE1.hal1 0.002 0.044 0.044 93
TMLE2.hal1 0.000 0.034 0.034 94
DR1.gbm5 0.012 0.044 0.043 92
DR2.gbm5 0.010 0.034 0.033 91
TMLE1.gbm5 0.009 0.044 0.043 93
TMLE2.gbm5 0.008 0.034 0.033 92
DR1.gbm1 0.010 0.044 0.043 92
DR2.gbm1 0.009 0.034 0.033 92
TMLE1.gbm1 0.007 0.044 0.043 92
TMLE2.gbm1 0.006 0.034 0.033 92
Table 1: Bias, empirical standard error, mean of standard error estimates, and percentage coverage of 95% confidence intervals for Scenario 1.
bias empSE SEhat cover
HT 0.002 0.046 0.046 94
Haj 0.000 0.037 0.037 95
naive -0.288 0.035
DR1 0.001 0.052 0.051 96
DR2clw -0.001 0.044 0.043 95
DR2 -0.001 0.044 0.043 95
TMLE1 0.001 0.052 0.051 96
TMLE2 -0.001 0.044 0.043 95
KH 0.001 0.052 0.051 96
DR1.hal5 0.002 0.052 0.049 95
DR2.hal5 0.001 0.044 0.041 94
TMLE1.hal5 0.002 0.052 0.050 95
TMLE2.hal5 0.000 0.044 0.042 94
DR1.hal1 0.002 0.052 0.050 95
DR2.hal1 0.001 0.044 0.041 94
TMLE1.hal1 0.001 0.052 0.050 95
TMLE2.hal1 0.000 0.044 0.042 94
DR1.gbm5 0.021 0.058 0.053 91
DR2.gbm5 0.019 0.051 0.045 90
TMLE1.gbm5 0.020 0.052 0.054 92
TMLE2.gbm5 0.019 0.044 0.045 91
DR1.gbm1 0.017 0.052 0.048 92
DR2.gbm1 0.016 0.043 0.039 90
TMLE1.gbm1 0.013 0.052 0.049 94
TMLE2.gbm1 0.011 0.044 0.040 92
Table 2: Bias, empirical standard error, mean of standard error estimates, and percentage coverage of 95% confidence intervals for Scenario 2.
bias empSE SEhat cover
HT 0.001 0.062 0.063 94
Haj -0.001 0.053 0.054 95
naive -0.399 0.047
DR1 0.222 0.068 0.078 16
DR2clw 0.184 0.053 0.069 18
DR2 0.220 0.058 0.069 8
TMLE1 0.125 0.064 0.070 58
TMLE2 0.124 0.053 0.060 46
KH 0.089 0.062 0.063 72
DR1.hal5 0.002 0.063 0.062 94
DR2.hal5 0.001 0.053 0.053 94
TMLE1.hal5 0.002 0.063 0.066 96
TMLE2.hal5 0.001 0.053 0.053 94
DR1.hal1 0.002 0.062 0.062 94
DR2.hal1 0.001 0.053 0.053 94
TMLE1.hal1 0.002 0.062 0.067 96
TMLE2.hal1 0.001 0.053 0.053 94
DR1.gbm5 0.020 0.063 0.062 94
DR2.gbm5 0.018 0.053 0.052 92
TMLE1.gbm5 0.016 0.063 0.066 96
TMLE2.gbm5 0.015 0.053 0.052 92
DR1.gbm1 0.019 0.063 0.062 94
DR2.gbm1 0.018 0.053 0.052 92
TMLE1.gbm1 0.016 0.063 0.066 95
TMLE2.gbm1 0.014 0.053 0.052 92
Table 3: Bias, empirical standard error, mean of standard error estimates, and percentage coverage of 95% confidence intervals for Scenario 3.
bias empSE SEhat cover
HT -0.002 0.109 0.110 94
Haj -0.001 0.091 0.094 95
naive -0.399 0.081
DR1 0.238 0.119 0.126 56
DR2clw 0.190 0.084 0.110 60
DR2 0.239 0.099 0.110 42
TMLE1 0.118 0.110 0.116 86
TMLE2 0.119 0.090 0.098 82
KH 0.085 0.108 0.106 87
DR1.hal5 0.000 0.109 0.106 93
DR2.hal5 0.001 0.090 0.089 95
TMLE1.hal5 -0.000 0.108 0.113 94
TMLE2.hal5 0.001 0.090 0.089 95
DR1.hal1 0.000 0.108 0.106 93
DR2.hal1 0.001 0.090 0.089 95
TMLE1.hal1 -0.000 0.108 0.113 94
TMLE2.hal1 0.001 0.090 0.089 95
DR1.gbm5 0.024 0.109 0.105 93
DR2.gbm5 0.025 0.090 0.088 93
TMLE1.gbm5 0.019 0.109 0.113 95
TMLE2.gbm5 0.020 0.090 0.088 94
DR1.gbm1 0.020 0.109 0.105 94
DR2.gbm1 0.021 0.090 0.088 94
TMLE1.gbm1 0.016 0.109 0.113 96
TMLE2.gbm1 0.016 0.090 0.088 94
Table 4: Bias, empirical standard error, mean of standard error estimates, and percentage coverage of 95% confidence intervals for Scenario 4.
bias empSE SEhat cover
HT -0.011 0.111 0.112 94
Haj -0.013 0.092 0.093 94
naive -0.406 0.081
DR1 0.239 0.129 0.123 52
DR2clw 0.186 0.090 0.104 57
DR2 0.237 0.111 0.104 38
TMLE1 0.117 0.108 0.115 85
TMLE2 0.115 0.087 0.094 79
KH 0.082 0.105 0.107 90
DR1.hal5 -0.003 0.105 0.106 95
DR2.hal5 -0.005 0.086 0.086 94
TMLE1.hal5 -0.003 0.105 0.113 96
TMLE2.hal5 -0.005 0.086 0.086 94
DR1.hal1 -0.003 0.105 0.106 95
DR2.hal1 -0.005 0.086 0.086 94
TMLE1.hal1 -0.003 0.105 0.113 96
TMLE2.hal1 -0.005 0.086 0.086 94
DR1.gbm5 0.022 0.106 0.106 96
DR2.gbm5 0.021 0.087 0.085 94
TMLE1.gbm5 0.018 0.106 0.112 97
TMLE2.gbm5 0.016 0.087 0.086 94
DR1.gbm1 0.019 0.105 0.106 96
DR2.gbm1 0.017 0.086 0.085 94
TMLE1.gbm1 0.015 0.106 0.112 97
TMLE2.gbm1 0.013 0.087 0.086 94
Table 5: Bias, empirical standard error, mean of standard error estimates, and percentage coverage of 95% confidence intervals for Scenario 5.
bias empSE SEhat cover
HT 0.008 0.195 0.196 95
Haj 0.004 0.164 0.164 95
naive -0.394 0.142
DR1 0.329 0.262 0.212 68
DR2clw 0.224 0.147 0.178 80
DR2 0.325 0.246 0.178 58
TMLE1 0.123 0.186 0.202 94
TMLE2 0.119 0.149 0.165 91
KH 0.095 0.178 0.184 94
DR1.hal5 0.010 0.180 0.184 95
DR2.hal5 0.007 0.146 0.149 96
TMLE1.hal5 0.011 0.180 0.196 97
TMLE2.hal5 0.007 0.145 0.149 96
DR1.hal1 0.011 0.180 0.184 95
DR2.hal1 0.007 0.145 0.149 96
TMLE1.hal1 0.010 0.180 0.196 97
TMLE2.hal1 0.006 0.146 0.149 96
DR1.gbm5 0.035 0.180 0.184 94
DR2.gbm5 0.031 0.146 0.149 95
TMLE1.gbm5 0.036 0.181 0.196 96
TMLE2.gbm5 0.032 0.146 0.149 94
DR1.gbm1 0.032 0.181 0.184 95
DR2.gbm1 0.029 0.146 0.149 95
TMLE1.gbm1 0.032 0.181 0.196 96
TMLE2.gbm1 0.028 0.146 0.149 95
Table 6: Bias, empirical standard error, mean of standard error estimates, and percentage coverage of 95% confidence intervals for Scenario 6.

### v2

@@ 6 Ratio estimator
This may be more efficient than $\hat{\theta}_{1}$, for the same reason that the Hajek estimator may be more efficient than the Horwitz-Thompson estimator. In CLW’s simulation study with parametric nuisance models, $\hat{\theta}_{{\rm CLW}2}$ did indeed have smaller variance than $\hat{\theta}_{1}$.
@@
The mean-zero term $\sqrt{M}n^{-1}\sum_{i=1}^{n}\{m_{0}(X_{i})-Y_{i}\}$ in equation (7) is $O_{p}(1)$ and does not vanish asymptotically. However, when the population is large compared to Samples A and B, its variance is small compared to that of the other terms in expression (7). We see that if this term is ignored, equation (7) is the same as equation (4) but with $m_{0}(X_{i})$ in the first term replaced by $m_{0}(X_{i})-n^{-1}\sum_{i^{\prime}=1}^{n}m_{0}(X_{i^{\prime}})$.
The asymptotic variance of $\hat{\theta}_{{\rm CLW}2}-\bar{Y}$ (and hence of $\hat{\theta}_{2}-\bar{Y}$) is given by CLW (who also ignored the $\sqrt{M}n^{-1}\sum_{i=1}^{n}\{m_{0}(X_{i})-Y_{i}\}$ term), along with an estimator of this variance. This is described in more detail by Seaman et al. (2025)[26] (specifically, the formulae for $\mu_{\rm DR2}$ when both nuisance models are correctly specified).
@@ 7 Targeted maximum likelihood
A popular alternative to estimator $\hat{\theta}_{\rm iid}$ for iid data is the targeted maximum likelihood estimator (TMLE). The TMLE method involves modifying an initial estimate $\hat{m}_{\rm iid}$ of $E(Y\mid X)$ to $\hat{m}^{*}_{\rm iid}$ in such a way that the simple regression imputation estimator $n^{-1}\sum_{i=1}^{n}\hat{m}_{\rm iid}^{*}(X_{i})$ of $E(Y)$ can be validly used even if $\hat{m}_{\rm iid}$ was obtained using a data-adaptive method. We now describe an analogous TMLE method for our survey sample situation.
@@
for some $\epsilon_{k}$. Alternatively, if $Y$ is bounded by zero and one, i.e. $P(0<Y<1)=1$, then $\hat{m}_{k}(x;\epsilon_{k})$ can be defined by either equation (8) or equation (9):
@@ 8 Calculating $\hat{\pi}^{B}_{k}$ and $\hat{m}_{k}$
So far, we have been agnostic about how to calculate $\hat{\pi}^{B}_{k}$ and $\hat{m}_{k}$, except to specify that they only use, respectively, data $\{(R^{A}_{i},R^{B}_{i},R^{A}_{i}\pi^{A}_{i},R^{A}_{i}X_{i},R^{B}_{i}X_{i}):i\notin\mathcal{S}_{k}\}$ and data $\{(R^{B}_{i},R^{B}_{i}X_{i},R^{B}_{i}Y_{i}):i\notin\mathcal{S}_{k}\}$ not in fold $k$ (and, when calculating $\hat{\pi}^{B}$, setting $R^{A}=R^{A}X=0$ for individuals not in active clusters and compensating for this by scaling the $\pi^{A}_{i}$’s). Estimating $m_{0}(X)=E(Y\mid X)$ is relatively straightforward, because $Y_{1},\ldots,Y_{n}$ are assumed to be independent given $X_{1},\ldots,X_{n}$. This is a standard problem for which many data-adaptive methods could be used. In the simulation study of Section 9 we use both XGBoost and Highly Adaptive LASSO (HAL).
@@
This is the log likelihood for a binomial regression model of $R^{B}$ on $X$ in the population of $n$ individuals, using the individuals in Sample A weighted by their inverse sampling probabilities to represent the population. Ferri-Garcia et al. (2022)[16] (see also [11]) use the pseudo-likelihood together with a classification and regression tree to estimate $\pi^{B}_{0}(X)$. In Section 9, we use expression (11) as an objective function (to be maximised) for XGBoost.
The approximate log pseudo-likelihood is the same as expression (11) but omitting the $R^{B}_{i}\log\{1-\pi^{B}(X_{i})\}$ term. When Sample B is small compared to the population, this term will be negligible compared to the other terms in expression (11). The approximate pseudo-likelihood has the advantage that it can be used quite generally with data-adaptive methods that predict a binary outcome and allow for weights. In Section 9, we use the approximate pseudo-likelihood with HAL. We acknowledge CLW’s criticism of the approximate pseudo-likelihood in situations where Sample B is not a small fraction of the population, but we focus in Section 9 on scenarios where Sample B is small compared to the population, a situation which would be quite common in practice.
@@ 9 Simulation study
We simulated a finite population with $J=1000$ clusters. Each cluster $j$ contains $H_{jq}$ households of $q$ individuals ($q=1,2,3$), where $H_{jq}$ is negatively binomially distributed with mean 100 and variance 400. Hence, the expected number of individuals in a cluster is 600, and the expected population size is 600,000. For each individual $i$ in cluster $j$, continuous variables $X_{1i}$ and $X_{2i}$ and binary variables $X_{3i}$ and $X_{4i}$ were generated independently, with expectations that depend on the total number of households in the cluster $j$, the size (1, 2 or 3) of the household to which individual $i$ belongs, and a cluster-level random effect. See Appendix A9 for details. This fixed population was used for all the simulations.
For each simulated dataset, $Y_{1},\ldots,Y_{n}$ were generated independently from a normal distribution with mean that depends on $X=(X_{1},X_{2},X_{3},X_{4})$ and variance 1. Sample A was drawn using a three-stage sampling design. First $M$ clusters were selected using Sampford sampling, with the probability of selecting cluster $j$ being proportional to the total number of households in cluster $j$. From each of these $M$ selected clusters, $n_{\rm house}$ households were selected using SRSWOR. Then one individual was sampled at random from each of these $n_{\rm house}$ sampled households. Thus, Sample A contained $M\times n_{\rm house}$ individuals.
We considered 24 scenarios. In Scenarios 1–12, $m_{0}(X)$ is linear in $X$ and $\pi^{B}_{0}(X)$ follows a logistic regression model with only main effects. Specifically, $m_{0}(X)=X_{1}+X_{2}+2X_{3}+X_{4}$ and $\mbox{logit }\pi^{B}_{0}(X)=\alpha_{\rm int}+0.5X_{1}+X_{2}+0.5X_{3}+X_{4}$. In Scenarios 13–24, $m_{0}(X)$ and $\pi^{B}_{0}(X)$ each include an interaction and a quadratic term. Specifically, $m_{0}(X)=0.5X_{1}+0.5X_{2}+2X_{3}+X_{4}+2X_{1}X_{3}+X_{2}^{2}$ and $\mbox{logit }\pi^{B}_{0}(X)=\alpha_{\rm int}+0.25X_{1}+0.5X_{2}+0.5X_{3}+X_{4}+X_{1}X_{3}+0.5X_{2}^{2}$. The value of $\alpha_{\rm int}$ was chosen to achieve a desired expected size $E(n^{B})$ of Sample B. Each of these two sets of 12 scenarios corresponded to two choices of $M$ (150 and 50), two choices of $n_{\rm house}$ (20 and 5), and three choices of $E(n^{B})$ (5000, 1000 and 500).
The following estimators were applied to each of 400 simulated datasets for each of the 24 scenarios.
@@
: $\hat{\theta}_{1}$ using HAL to estimate $\pi^{B}_{0}$ and $m_{0}$ with cross-fitting
@@
: $\hat{\theta}_{2}$ using HAL to estimate $\pi^{B}_{0}$ and $m_{0}$ with cross-fitting
@@
: $\hat{\theta}_{\rm TMLE1}$ using HAL to estimate $\pi^{B}_{0}$ and $m_{0}$ with cross-fitting
@@
: $\hat{\theta}_{\rm TMLE2}$ using HAL to estimate $\pi^{B}_{0}$ and $m_{0}$ with cross-fitting
Of these, DR1, DR2clw and KH were described by CLW. The parametric nuisance models we used were linear regression (for $m_{0}$) and logistic regression (for $\pi^{B}_{0}$) models with only main effects for $X_{1}$, $X_{2}$, $X_{3}$ and $X_{4}$. Hence, these models are correctly specified in Scenarios 1–12 and misspecified in Scenarios 13–24. For DR1, DR2clw, TMLE1 and TMLE2, we estimated the standard error (SE) and calculated 95% confidence intervals (CIs) using variance estimators that are valid when the model for $\pi^{B}_{0}$ (but not necessarily $m_{0}$) is correctly specified[26]. In addition, we calculated $\hat{\theta}_{1}$, $\hat{\theta}_{2}$, $\hat{\theta}_{\rm TMLE1}$ and $\hat{\theta}_{\rm TMLE2}$ using XGBoost, instead of HAL, to estimate $\hat{\pi}^{B}$ and $\hat{m}$. These will be denoted ‘DR1.xgb5’ and so on. For methods using cross-fitting, we used $K=5$ folds, set $\delta=0.01$ and used $L=4$ sets of clusters based on quartiles of the distribution of $\pi^{C}_{j}$ (as described at the end of Section 5). We also applied the estimators that used HAL or XGBoost without cross-fitting (denoted ‘DR1.hal1’, ‘DR1.xgb1’, and so on).
Full results are given in Appendix A10. Here, we highlight the main findings. In Scenarios 1–4, where $E(n^{B})=5000$, all estimators except Naive were approximately unbiased with CIs that had approximately correct coverage. The estimators that used HAL or XGBoost with cross-fitting to estimate the nuisance functions had very similar SE to the corresponding estimators that used parametric nuisance models (see Figure 1).
Figure 1: Empirical standard error of $\hat{\theta}_{1}$ when $\pi^{B}_{0}$ and $m_{0}$ are estimated using parametric models (DR1), HAL (DR1.hal5) or XGBoost (DR1.xgb5), for Scenarios 1–12.
The Hajek, DR2 and TMLE2 estimators were more efficient than the corresponding HT, DR1 and TMLE1 estimators; the SEs of the former were approximately half that of the former. Findings for Scenarios 5–8, where $E(n^{B})=1000$, were similar, although CI coverage of DR2.xgb5 and TMLE2.xgb5 dropped to 89% in Scenario 6, where $M=150$ and $n_{\rm house}=5$. Coverage for DR2.hal5, TMLE2.hal5, DR2.xgb5 and TMLE2.xgb5 dropped further in Scenarios 9–12, where $E(n^{B})=500$, but usually remained above 90% and was never less than 87%.
In Scenarios 13–24, the estimators using parametric nuisance models were biased and had CIs with undercoverage (see Figure 2).
Figure 2: Left: Coverage of 95% confidence interval for $\bar{Y}$ when $\hat{\theta}_{1}$ used with parametric nuisance models (DR1), HAL (DR1.hal5), XGBoost (DR1.xgb5) or Kim and Haziza’s method (KH), for Scenarios 13–24. Right: Coverage when $\hat{\theta}_{{\rm CLW}2}$ used with parametric nuisance models (DR2clw) or $\hat{\theta}_{2}$ used with HAL (DR2.hal5) or XGBoost (DR2.xgb5).
Coverage was particularly bad (in one case, as low as 11%) for the DR2clw and TMLE2 estimators. Estimators using HAL or XGBoost had much smaller bias, especially in Scenarios 13–16, where $E(n^{B})=5000$; here bias was almost zero. CI coverage for these estimators was approximately 95% and never dropped below 90%.
Using HAL or XGBoost without cross-fitting appeared to have little effect on bias and virtually no effect on the empirical SE. For XGBoost (but less obviously for HAL), omitting the cross-fitting step caused CI coverage to deteriorate in many of the scenarios, especially for DR2 and TMLE2 (see Figure 3). This lack of obvious effect on bias and this deterioration of coverage is in line with the findings of Ellul et al. (2025)[15].
Figure 3: Coverage of 95% confidence interval for $\bar{Y}$ when $\hat{\theta}_{1}$ (top) or $\hat{\theta}_{2}$ (bottom) is used with HAL (left) or XGBoost (right). Each plot compares coverage when 5-fold cross-fitting is used versus when no cross-fitting is used. Each dot corresponds to one of the 24 scenarios. Broken lines indicate perfect 95% coverage.
Finally, we also calculated $\hat{\theta}_{2}$ using parametric nuisance models. As expected, the biases and SEs of this estimator were very similar to those of DR2clw. Over the 24 scenarios, the mean (maximum) absolute difference in their biases was 0.0011 (0.0036), and the ratio of their SEs varied from 0.99 to 1.01.